A heat engine takes in 360J of energy from a hot reservoir and performs 25.0J of work in each cycle. Calculate:
(a) the efficiency of the engine
(b) the energy expelled to the cold reservoir in each cycle.
1
Expert's answer
2021-03-01T11:51:28-0500

"\eta=\frac{Q_2}{Q_1}=\frac{25}{360}\approx 0.07=7\%."

"\Delta Q=Q_1-Q_2=360-25=335~J."